OpenStudy (anonymous):

solve the following quadratic function x^2-16x+63?

OpenStudy (anonymous):

\[X=( -b \pm \sqrt{a^2 -4ac} )\div 2c\]

OpenStudy (anonymous):

(x - 7) • (x - 9)

OpenStudy (anonymous):

ax2+bx+c

OpenStudy (anonymous):

Step 1 : Simplify x2-16x + 63 Trying to factor by splitting the middle term 1.1 Factoring x2-16x+63 The first term is, x2 its coefficient is 1 The middle term is, -16x its coefficient is -16 The last term, "the constant", is +63 Step-1 : Multiply the coefficient of the first term by the constant 1 • 63 = 63 Step-2 : Find two factors of 63 whose sum equals the coefficient of the middle term, which is -16 -63 + -1 = -64 -21 + -3 = -24 -9 + -7 = -16 That's it Step-3 : Rewrite the polynomial splitting the middle term using the two factors found in step 2 above, (-9) and (-7) x2 - 9x - 7x - 63 Step-4 : Add up the first 2 terms, pulling out like factors : x • (x-9) Add up the last 2 terms, pulling out common factors : 7 • (x-9) Now add up the four terms of step 3 : (x-7) • (x-9) Which is the desired factorization (x - 7) • (x - 9)

OpenStudy (jojo4eva):

(x-7) (x-9)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome!

OpenStudy (anonymous):

:)